find the sum of first 99 natural no.s.(arithmetic progression)

Grade:10

1 Answers

Abu Azbar
6 years ago
n=99Therefore : S = n/2(2a+(n-1)d) = 99/2(2*1+(91-1)1) = 99/2 *92 = 99*46 = 4554Therefore the sum of first 99 natural numbers is 4554